# Rethinking LLM-Based Recommendations: A Query Generation-Based, Training-Free Approach

### v1

@@ 4 Experiment
4.1 Datasets and Metrics
Dataset #Users #Items #Reviews Density (%)
Sports 35,598 18,357 296,337 0.0453
Beauty 22,363 12,101 198,502 0.0734
Toys 19,412 11,924 167,597 0.0724
Table 1: Statistics of the datasets.
We conducted experiments on three widely used benchmark datasets collected from the Amazon e-commerce platform: Sports & Outdoors, Beauty, and Toys & Games. We selected the three aforementioned domains for experimentation. Each dataset consists of user interactions, including a user ID, an item ID, a rating, a review, and a timestamp. The statistics of the data set are provided in Table 1. Following previous studies Li et al. (2023); Wang et al. (2024a), we partitioned the data in the same way that the last interaction of a user served as the test label, the second to last as the validation label, and the remaining interactions as training data.
To evaluate the performance of the model, we adopted two widely used ranking metrics: Hit Rate (HR) and Normalized Discounted Cumulative Gain (NDCG), with $k\in\{5,10\}$.
@@ 4.1 Datasets and Metrics
Table 2: Evaluation results in the RecSys-LLM cooperation scenario. The best and second-best results for each metric are highlighted in bold and underlined, respectively. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
4.2 Experimental Setup and Baselines
To evaluate the effectiveness of QueRec, we integrated it into an existing CF-based recommender system and compared its performance against existing LLM-based reranking methods. Our goal was to examine whether the proposed parallel architecture would be more effective than the traditional serial reranking approach. For scalability, we employed LLaMA3.2 3B Dubey et al. (2024) for query generation and the implementation of LLM-based baselines, while for similarity-based retrieval, we leveraged a widely used pre-trained encoder Li and Li (2023) without any additional training.
In the first experiment, we evaluated the performance of QueRec in comparison to existing LLM-based reranking methods when applied to a representative sequential recommender system. We selected SASRec Kang and McAuley (2018), DIF-SR Xie et al. (2022), and ELMRec Wang et al. (2024a) as baselines. The existing LLM-based methods perform a reranking of the top 20 items selected by these CF-based models. LLM-based reranking methods include three variations of the LLMRank Hou et al. (2024) without training: sequential, in-context learning, and recency-focused, as well as TALLRec Bao et al. (2023) in two configurations: zero shot and fine-tuned.
In the second experiment, we compared QueRec as a standalone method against traditional training-based approaches. To better reflect real-world recommendation scenarios with a large candidate pool, we conducted all experiments without separate negative item selection, evaluating performance over the entire item pool.
More details on the experimental settings and baselines are provided in Appendix and in the online repository [^1] https://anonymous.4open.science/r/querec/.
@@ 4.3 Main Results
Model Sports Beauty Toys
@@
QueRec 0.0280 0.0188 0.0403 0.0228 0.0501 0.0354 0.0675 0.0410 0.0680 0.0472 0.0937 0.0555
@@
Table 3: Performance comparison of sequential recommendation models. Metrics with lower performance than QueRec are shaded in blue, while higher-performing cases are shaded in red.
The first experiment evaluated the effectiveness of integrating the proposed method into an existing recommendation system to enhance its performance. The results of this experiment are presented in Table 2. We observed that previously proposed LLM-based reranking methods lead to a decrease in performance. This decline seems to be due to the fact that the top 20 candidate items provided by the CF-based model are very similar to each other, which may confuse the LLM during reranking. Additionally, in the case of TALLRec, performance deteriorated in most scenarios when fine-tuning was applied. These results suggest that learning CF-signals solely through natural language-based training remains a challenging task, even when the LLM is trained on candidate items from CF-based models.
In contrast, QueRec achieved an average performance improvement of 31% and a maximum improvement of 57%. In particular, QueRec does not require additional training when integrated into an existing recommendation system, highlighting its efficiency and scalability. These results demonstrate that QueRec can effectively leverage the world knowledge inherent in LLM, thereby providing differentiated recommendations from CF-based models.
@@
In the second experiment, we compared the performance of QueRec as a standalone method against traditional sequential recommendation models. The results presented in Table 3 demonstrate that QueRec achieves competitive zero shot performance without additional training, outperforming six of the eight trained traditional recommendation models. This finding is particularly notable given that previous studies have indicated that zero-shot LLM-based approaches have failed to surpass full-shot models. These results suggest that QueRec effectively leverages the knowledge and reasoning capabilities of LLMs, providing a significant advantage in generating queries compared to existing reranking methods.
In addition, we conducted an ablation study to assess the contribution of each component in our approach. As shown in Table 3, we evaluated performance changes when each of the following three components was removed: (1) recent item information, (2) item queries, and (3) user queries. Among these, the absence of user queries resulted in the most significant performance degradation, while excluding the other components also led to notable declines. These findings indicate that all proposed components contribute to the improvement in performance, confirming the effectiveness of QueRec.
4.4 In-depth Analysis
In this subsection, we analyze the distinguishing characteristics and advantages of the proposed method. Since QueRec relies solely on pre-trained models without additional training, it effectively leverages the diverse knowledge embedded in LLMs while avoiding bias toward any specific domain.
@@ 4.4.1 Diversity of Recommendations
Our proposed method does not rely on domain-specific training, allowing unbiased recommendations, which in turn leads to more diverse item suggestions. Traditional recommendation systems learn not only collaborative signals, but also item popularity, through training. Although this can improve overall performance metrics, it can also lead to unwanted side effects, such as filter bubbles, where frequently recommended items continue to dominate recommendations Nguyen et al. (2014). To evaluate whether the proposed method improves the diversity of recommendations, we compare the distribution of recommended item IDs with those from conventional recommendation systems.
Fig. 3 illustrates the distribution of recommended items between QueRec and the baselines. Both SASRec and DIF-SR exhibit a bias toward lower-ID item groups, suggesting that the models amplify the bias present in the dataset through the training process. In contrast, ELMRec demonstrates distinct distributions, yet still exhibits skewness toward specific items. This bias is likely a result of its multitask learning in a text-to-text format, but it still struggles to maintain recommendation diversity. In contrast, our proposed method, QueRec, shows a more balanced distribution of recommended items, indicating a stronger trend towards diverse recommendations. This characteristic has practical implications for real-world recommendation systems, as it can mitigate filter bubbles and facilitate a more exploratory recommendation experience based on user preferences. These findings highlight that QueRec not only incorporates a distinct perspective compared to conventional recommendation models, but also has the potential to generate synergies when integrated with existing approaches. A more detailed analysis of this can be found in the Appendix B.2.
4.4.2 Performance in Minor Items
Figure 4: Performance comparison on minor item group in three datasets.
To examine whether the observed improvement in the diversity of recommendations improves the performance of minor items, we performed an analysis using the 20% of items that appeared less frequently in the training set as the minor group. Fig. 4 presents the performance of the recommendation of each method for these elements. In most cases, our proposed method outperformed existing approaches, demonstrating its effectiveness in improving recommendations for items observed less frequently.
However, in the Sports dataset, QueRec exhibited performance comparable to that of ELMRec, likely due to ELMRec’s exposure to a broader range of tasks during training. Given that QueRec does not require additional training, these results offer valuable insights into how our method can mitigate the limitations of traditional recommendation systems while leveraging LLMs’ inherent knowledge and generalization capabilities.
@@ 5 Conclusion
In this work, we introduced QueRec, a novel framework that enhances the diversity and precision of recommendations by using LLM to generate feature-aware queries. Unlike LLM-based reranking approaches rely on pre-selected candidate sets, QueRec enables direct retrieval from the entire item pool, mitigating performance bottlenecks while achieving competitive zero-shot performance that often exceeding trained baselines. Furthermore, our analysis highlights the ability of the method to reduce bias, improve recommendations for minor items, and ultimately foster a more diverse recommendation. By integrating the complementary strengths of LLMs and recommendation models, QueRec provides a promising direction for future recommendation research, offering a practical, training-free solution adaptable to evolving LLMs and large-scale dynamic item pools.
@@ 6 Limitations
Despite its advantages, QueRec has several limitations. First, while it eliminates the need for explicit candidate selection, the effectiveness of query generation depends on the quality of the pre-trained LLM and its ability to generalize across domains. In cases where the LLM lacks sufficient domain-specific knowledge, the generated queries may not capture user intent effectively. Second, since our approach relies on retrieving items based on semantic similarity, it may underperform in domains where structured interaction data (e.g., collaborative filtering signals) are more informative than textual item representations. Finally, while QueRec enhances diversity, it does not explicitly control for fairness, popularity bias, or serendipity, which remain open challenges for future research. Addressing these limitations through hybrid approaches, enhanced query generation mechanisms, and domain-aware enhancements presents a promising direction for further improving LLM-driven recommendation systems.
@@ 7 Ethics Statement
Our research leveraged open source tools, including PyTorch Paszke et al. (2019), along with pre-trained language models such as LLaMA3.2 and T5 obtained via the Huggingface Wolf et al. (2019) library.
@@
Appendix A Baselines
A.1 Recommendation Model
@@
  • P5 Geng et al. (2022): A T5-base text-to-text recommendation model, converting all recommendation-related data into natural language sequences and utilizing personalized prompts for various recommendation tasks.
To obtain similarity scores for SASRec and DIF-SR, we trained the models using Recbole [^2] https://recbole.io/, a well-validated open-source recommendation framework. Detailed hyperparameters and training results will be made available through our public repository. For ELMRec, we excluded test targets affected by label leakage in the explanation task to ensure fair evaluation. In case of ELMRec, we obtain beam search scores and use them as the recommendation score. The model was trained using the same T5-small backbone and hyperparameters as in prior studies. For all other baselines, we report the performance metrics as documented in previous studies Zhou et al. (2020); Li et al. (2023); Wang et al. (2024a).
A.2 LLM-based Reranking Method
@@
  • TALLRec Bao et al. (2023): An approach that fine-tunes LLMs using LoRA to enhance their recommendation capabilities, enabling efficient adaptation to recommendation tasks while reducing computational overhead. The initially proposed method used a binary prediction approach, whereas we modified the prompt to select the top-10 items from the candidate set used in a previous study Kim et al. (2024). During training, the prompt included the top-20 candidate items recommended by the traditional recommendation model, while the training target was a reordered list of 10 item titles, ensuring that the label item appeared first. This setup was designed to allow a fair comparison that closely aligns with our proposed method. We trained the LLaMA 3.2 3B with LoRA for 2 epochs with a batch size of 4 and a learning rate of 5e-5.
Appendix B Additional Experiments
In this section, we present additional experiments that were not included in the main text due to space constraints.
B.1 LLM Variation
To evaluate the generalizability of our proposed approach across different LLM architectures, we conducted experiments using various models, including LLaMA 3.1 8B, LLaMA 3.3 70B, Gemma 2 2B, and Gemma 2 3B. Table 4 presents the performance of QueRec in a standalone setting. For the 70B model, we utilized Llama-3.3-70B-Instruct-Turbo in DeepInfra [^3] https://deepinfra.com/ for generation, as it was not feasible to run on our locally available GPUs.

### v2

@@ 4 Experiment
This section presents the experimental results and analysis that demonstrate the effectiveness of our proposed method.
@@
Table 1: Evaluation results in the LLM-recommender cooperation scenario. The best and second-best results for each metric are highlighted in bold and underlined, respectively. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
4.1 Experimental Setup
To evaluate the effectiveness of QueRec, we integrate existing CF-based recommendation method into our method and compared its performance against existing LLM-based reranking methods. Our goal was to examine whether the proposed parallel architecture would be more effective than the traditional serial reranking approach. For scalability, we employed LLaMA3.2 3B Dubey et al. (2024) for query generation and the implementation of LLM-based baselines, while for similarity-based retrieval, we leveraged a widely used pre-trained encoder Li and Li (2023) without any additional training.
➊ In the first experiment, we evaluated the performance of QueRec in comparison to existing LLM-based reranking methods when adapting sequential recommendation models. For the existing LLM-based reranking methods, recommendation models are used as candidates selectors similar to the previous study Yang et al. (2023). ➋ In the second experiment, we compared QueRec against existing CF-based approaches, T5-based approaches, and retrieval-based approaches that involve fine-tuning LLMs for query generation.
To better reflect real-world recommendation scenarios with a large candidate pool, we evaluated performance over the entire item pool. We adopted two widely used ranking metrics: Hit Rate (HR) and Normalized Discounted Cumulative Gain (NDCG), with $k\in\{5,10\}$. More details on the experimental settings and implimentation are provided in the appendix and online repository [^1] https://github.com/venzino-han/QueRec2025.
@@ 4.2 Main Results
In this subsection, we present key results comparing our method with existing baselines.
4.2.1 LLM-recommender Cooperation Scenario
Method Model Sports Beauty Toys
@@ 4.2.1 LLM-recommender Cooperation Scenario
POD 0.0496 0.0396 0.0576 0.0419 0.0537 0.0395 0.0688 0.0443 0.0691 0.0599 0.0742 0.0610
RDRec 0.0505 0.0408 0.0596 0.0433 0.0601 0.0461 0.0743 0.0504 0.0723 0.0593 0.0802 0.0605
Retrieval-based GPT4Rec 0.0002 0.0001 0.0004 0.0002 0.0004 0.0002 0.0006 0.0003 0.0005 0.0003 0.0011 0.0005
BIGRec 0.0071 0.0046 0.0115 0.0061 0.0237 0.0169 0.0355 0.0207 0.0324 0.0229 0.0471 0.0276
Ours QueRec 0.0589 0.0455 0.0746 0.0506 0.0816 0.0625 0.1060 0.0704 0.1092 0.0777 0.1350 0.0860
Table 2: Performance comparison of existing recommendation methods. The best results for each metric are highlighted in bold. "H@k" and "N@k" denote Hit Rate and NDCG at rank k, respectively.
The first experiment evaluated the effectiveness of our proposed parallel architecture. The results of this experiment are presented in Table 1. We observed that existing LLM-based reranking methods lead to a decrease in performance. This decline seems to be due to the fact that the selected candidate items from the CF-based model are very similar to each other, which may confuse the LLM during reranking. Although A-LLMRec, which uses CF-based embeddings, outperformed other LLM-based reranking methods, it still underperformed compared to our method. These results suggest that learning CF-signals through natural language-based training remains a challenging task, even when the LLM is trained on candidate items from CF-based models.
In contrast, QueRec achieved an average performance improvement of 31% and a maximum improvement of 57%. In particular, QueRec does not require additional training when integrated into an existing recommendation system, highlighting its epandability and ease of deployment. These results demonstrate that QueRec can effectively leverage the knowledge inherent in LLM, thereby providing differentiated recommendations from CF-based models.
4.2.2 Compare to Traditional Recommenders
In the second experiment, we compared the performance of QueRec against traditional sequential recommendation models and LLM-based retrieval methods. The results presented in Table 2 demonstrate that QueRec outperforms existing recommendation models. Furthermore, QueRec significantly outperformed existing LLM fine-tuning methods such as GPT4Rec and BIGRec, which generate item-titles and use them as queries. These results suggest that approaches based on fine-tuning may introduce biases induced by datasets and potentially compromise the inherent generalization capabilities of LLMs.
Method Sports Beauty Toys
H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10 H@5 N@5 H@10 N@10
QueRec 0.0589 0.0455 0.0746 0.0506 0.0816 0.0625 0.1060 0.0704 0.1092 0.0777 0.1350 0.0860
w/o CF-based model 0.0280 0.0188 0.0403 0.0228 0.0501 0.0354 0.0675 0.0410 0.0680 0.0472 0.0937 0.0555
@@ 4.2.2 Compare to Traditional Recommenders
Table 3: Evaluation results of ablation study. The best results for each metric are highlighted in bold.
4.3 In-depth Analysis
In this subsection, we analyze the distinguishing advantages of the proposed method.
@@ 4.3 In-depth Analysis
4.3.1 Ablation Study
We conducted an ablation study to assess the contribution of each component in our approach. As shown in Table 3, we evaluated performance changes when each of the key components was removed: (1) CF-based model (2) recent item information, (3) generated item description, and (4) user queries. The absence of user queries resulted in the most significant performance degradation, while excluding the other components also led to notable declines. Another notable finding is that our method outperforms fine-tuned LLM-based approaches (GPT4Rec and BIGRec), even without being combined with a CF-based model. This result is encouraging, as our method does not require any LLM fine-tuning steps.
@@ 4.3.2 Diversity of Recommendations
Our proposed method does not rely on additional training, allowing unbiased and more diverse recommendations compared to traditional methods that often reinforce popularity bias through training. As shown in Fig.3, existing methods exhibit skewed item distributions, suggesting that the training process introduces biases into these models. In contrast, our proposed method, produces a more balanced distribution of recommendation, indicating improved diversity.
This characteristic can mitigates filter bubbles and promotes exploratory user experiences. These findings highlight that our method not only incorporates a distinct perspective compared to traditional recommendation models, but also has the potential to generate synergies when integrated with existing approaches. Further analysis is provided in AppendixE.1.
Figure 4: Query quality comparison in the Toys Dataset. Related attributes are highlighted in the same color.
Figure 5: Novelty evaluation in three datasets.
4.3.3 Novelty Evaluation
To examine whether the observed improvement in the diversity of recommendations translates into the recommendation of more novel items, we analyze based on the average novelty scores of items in the Hit@10 results. Novelty is computed as the negative logarithm of the relative item popularity Zhou et al. (2010), defined as:
$\text{Novelty}(i)=-\log\left(\frac{\text{freq}(i)}{\sum_{j\in\mathcal{I}}\text{freq}(j)}\right),$ (11)
where $\text{freq}(i)$ denotes the frequency of item $i$ in the training set, and $\mathcal{I}$ is the set of all items.
Fig. 5 presents the average novelty scores of the top-10 recommended items for each method. In most cases, our proposed method outperformed existing approaches, indicating its effectiveness in promoting the recommendation of less popular and more novel items. These results suggest that our proposed method is capable of capturing fundamentally different signals compared to traditional recommendation systems, indicating its potential to deliver personalized recommendations tailored to users with diverse preferences.
4.3.4 Query Quality Analysis
To examine whether the proposed method effectively leverages the broad knowledge encoded in LLMs, we qualitatively compare the generated queries. Existing approaches, which are fine-tuned to produce item-title-style queries, often exhibit dataset-induced biases and struggle to provide clear reasoning for recommendations. In contrast, our method incorporates user history and reviews into the query, enabling retrieval based on richer contextual information and offering interpretability in the form of implicit rationales.
Fig. 4 presents example queries generated by BIGRec and QueRec, along with the actual items purchased by users. While QueRec does not explicitly mention the exact item titles, its queries align with the purchased items in terms of target user groups and product characteristics. Notably, our method generates semantically relevant queries that reflect the preferences of the user. These findings indicate that LLMs can leverage their knowledge for accurate and unbiased recommendations without requiring fine-tuning. Additional examples are included in Appendix E.3.1.
@@ 5 Conclusion
In this work, we introduced QueRec, a novel parallel framework that enhances the diversity and performance of recommendation by using LLM to generate personalized queries. QueRec outperforms state-of-the-art baselines including LLM-based reranking, LLM-based next-item prediction, and CF-based methods, by effectively leveraging the rich knowledge embedded in LLMs. Furthermore, our analysis highlights the ability of the method to reduce bias and increase the diversity of recommendations. By integrating the complementary strengths of LLMs and recommendation models, QueRec provides a promising direction for future recommendation research, offering a practical, training-free solution adaptable to evolving LLMs and large-scale dynamic item pools.
@@ 6 Limitations
Despite its advantages, QueRec has several limitations. First, while it eliminates the need for explicit candidate selection, the effectiveness of query generation depends on the quality of the pre-trained LLM and its ability to generalize across domains. In cases where the LLM lacks sufficient domain-specific knowledge, the generated queries may not capture user intent effectively. Second, since our approach relies on retrieving items based on semantic similarity, it may underperform in domains where structured interaction data (e.g., collaborative signals) are more informative than textual item representations. Finally, while QueRec enhances diversity, it does not explicitly control for fairness, popularity bias, or serendipity, which remain open challenges for future research. Addressing these limitations through hybrid approaches, enhanced query generation mechanisms, and domain-aware enhancements presents a promising direction for further improving LLM-driven recommendation systems.
@@ 7 Ethics Statement
Our research leveraged open source tools, including PyTorch Paszke et al. (2019), along with pre-trained language models such as LLaMA3.2 and T5 obtained via the Huggingface Wolf et al. (2019) library. We used all artifacts in accordance with their intended purpose.
@@
8 Acknowledgments
This research was supported by the National Research Foundation of Korea (RS-2025-16071337). For GPU infrastructure, our work was supported by the IITP grant funded by MSIT (No. RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale).
Appendix A Clarification of contribution
In this section, we further elaborate on the aspects that distinguish our work from prior studies and clarify the key contributions of this paper.
Early research on LLM-based recommendation systems primarily relied on fine-tuning LLMs as recommendation models Li et al. (2023a); Chen et al. (2024); Rajput et al. (2023); Geng et al. (2022). These methods either trained LLMs to generate user/item embeddings or directly predict item IDs. However, they often underutilize LLMs’ reasoning abilities and incur substantial training costs due to full-parameter updates. Another recent studies have proposed injecting LLM-derived user and item information into recommendation models to leverage LLM knowledge without fine-tuning the LLMs Xi et al. (2024); Zhang et al. (2025b). However, these methods require retraining the recommendation model to align with the LLM outputs. In contrast, we propose a training-free framework that fully leverages the pretrained LLM’s reasoning capacity. This design eliminates the need for fine-tuning, reducing bias and improving generalization.
Prior work such as GPT4Rec and BIGRec Li et al. (2023b); Bao et al. (2025) fine-tuned LLMs to generate the title-based queries. This restricts the expressiveness of the LLM and limits adaptability in dynamic item pools. Our method, by contrast, generates rich, personalized queries that incorporate user preferences and item attributes without training, enhancing retrieval quality and recommendation diversity.
Recent reranking approaches Hou et al. (2024); Bao et al. (2023); Kim et al. (2024); Yang et al. (2023) apply LLMs to reorder a small, pre-selected candidate set. However, these candidates are typically derived from CF-based models like SASRec Yang et al. (2023), constraining the LLM’s capacity to explore diverse knowledge. We address this limitation by using LLMs to generate both user and item queries for full-pool retrieval, enabling knowledge-rich, unbiased, and adaptive recommendations. Our proposed method also avoids the positional bias caused by the “lost in the middle” phenomenon, as it does not include candidate items in the prompt.
Another line of work involves using LLMs to generate synthetic queries, which are then used to train retrievers Mysore et al. (2023); Carranza et al. (2024). These studies have primarily been evaluated on search tasks where user intent is explicitly expressed through natural language queries. In contrast, our work focuses on recommendation settings, where user intent is implicit and not directly observable. Unlike prior efforts that use synthetic queries to train retrievers, we aim to leverage the knowledge encoded in LLMs to enrich limited user-item information and apply it to retrieval-based recommendation. Furthermore, we propose a parallel architecture that effectively integrates this augmented information with existing recommender models without requiring retraining.
Appendix B Experimental Setups
This section provides a more detailed description of the experimental setup and the baselines used.
B.1 Environment
For the query generation stage using the LLM, we employed the vLLM [^2] https://docs.vllm.ai/ framework, which enables efficient and scalable inference by leveraging optimized memory management and parallelization strategies.
B.2 Datasets
Dataset #Users #Items #Reviews Density (%)
Sports 35,598 18,357 296,337 0.0453
Beauty 22,363 12,101 198,502 0.0734
Toys 19,412 11,924 167,597 0.0724
Yelp 30,431 20,033 316,354 0.0519
Table 4: Statistics of the datasets.
We conducted experiments on four widely used benchmark datasets collected from the Amazon e-commerce platform: Sports & Outdoors, Beauty, Toys & Games and Yelp. Each dataset consists of user interactions, including a user ID, an item ID, a rating, a review, and a timestamp. The statistics of the data set are provided in Table 4. These datasets are widely adopted by previous studies Li et al. (2023c); Wang et al. (2024a, b); Rajput et al. (2023); Cao et al. (2024); Geng et al. (2022), which have been extensively explored over the past three years. Due to space limitations, we include only the experimental results for the three main datasets in the manuscript, excluding the Yelp dataset. To further validate the effectiveness of our proposed method, we conducted additional experiments on the Yelp dataset, which represents a domain distinct from the other three datasets. The results of experiment on the Yelp dataset are presented in Section F.
B.3 Baselines
This subsection introduces the baselines used in our experiments and their corresponding setups. To obtain similarity scores for SASRec and DIF-SR, we trained the models using Recbole [^3] https://recbole.io/, a well-validated open-source recommendation framework. Detailed hyperparameters and training results will be made available through our public repository. For ELMRec, we excluded test targets affected by label leakage in the explanation task to ensure fair evaluation. In case of ELMRec, we obtain beam search scores and use them as the recommendation score. The model was trained using the same T5-small backbone and hyperparameters as in prior studies. For all other baselines, we report the performance metrics as documented in previous studies Zhou et al. (2020); Li et al. (2023c); Wang et al. (2024a). In cases of fine-tuned LLM-based methods (e.g. TALLRec, BIGRec and GPT4Rec), we trained the LLaMA 3.2 3B with LoRA for 2 epochs with a batch size of 4 and a learning rate of 5e-5.
B.3.1 CF-base Model
@@
B.3.2 T5-base Model
  • P5 Geng et al. (2022): A T5-based text-to-text recommendation model, converting all recommendation-related data into natural language sequences and utilizing personalized prompts for various recommendation tasks.
  • POD Li et al. (2023c): A T5-based method, distills discrete prompts into continuous prompt vectors to better bridge user/item, aiming to improve both training efficiency and recommendation performance.
  • RDRec Wang et al. (2024b): A T5-based method, enhances recommendation by distilling rationales from user and item reviews, allowing a compact model to leverage preference and attribute-level explanations for improved recommendation performance.
@@ B.3.2 T5-base Model
B.3.3 LLM-based Reranking Method
@@
  • TALLRec Bao et al. (2023): An approach that fine-tunes LLMs using LoRA to enhance their recommendation capabilities, enabling efficient adaptation to recommendation tasks while reducing computational overhead. The initially proposed method used a binary prediction approach, whereas we modified the prompt to select the top-10 items from the candidate set used in a previous study Kim et al. (2024). During training, the prompt included the top-20 candidate items recommended by the traditional recommendation model, while the training target was a reordered list of 10 item titles, ensuring that the label item appeared first. This setup was designed to allow a fair comparison that closely aligns with our proposed method.
  • A-LLMRec Kim et al. (2024): An LLM-based framework that integrates pre-trained CF-based user/item embeddings into LLMs, enabling effective recommendations in both cold and warm scenarios. It leverages both collaborative signals and LLM reasoning capabilities by introducing a alignment layer that requires training to bridge the two representations.
B.3.4 LLM-based Retrieval Method
  • GPT4Rec Li et al. (2023b): An LLM-based recommendation method that fine-tunes language models to predict a user’s next item and uses the generated item text as search queries.
  • BIGRec Bao et al. (2025): A bi-step grounding method for LLM-based recommendation systems. It first aligns the LLM with the recommendation task by training it to generate item titles. These generated titles are then used as textual queries to retrieve items, thereby grounding the language model outputs to the recommendation space.
Appendix C Experiments on Various LLMs
To evaluate the generalizability of our proposed approach across different LLM architectures, we conducted experiments using various models, including LLaMA 3.1 8B, LLaMA 3.3 70B, Gemma 2 2B, and Gemma 2 3B. Table 5 presents the performance of QueRec in a standalone setting. For the 70B model, we utilized Llama-3.3-70B-Instruct-Turbo in DeepInfra [^4] https://deepinfra.com/ for generation.